OpenStudy (anonymous):

Please help! Solve -3x2 - 4x - 4 = 0. x equals quantity of 2 plus or minus 4i square root of 2 all over 3 x equals quantity of 2 plus or minus 2i square root of 2 all over 3 x equals quantity of negative 2 plus or minus 2i square root of 2 all over 3 x equals quantity of negative 2 plus or minus 4i square root of 2 all over 3

OpenStudy (mokeira):

\[-3x ^{2}-4x-4=0\] make sure x^2 is positive by multiplying everything by -1 to get \[3x ^{2}+4x+4=0\] Do you know how to factor out? @ibreezysaan

OpenStudy (anonymous):

No, sorry ): I'm not really good at it @Mokeira

OpenStudy (mokeira):

do you know quadratic formula?

OpenStudy (anonymous):

\[ -3x2 - 4x - 4 = 0\\ 3x^2+4x-4=0\] then the quadratic formula \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=3,b=4,c=-4\]

OpenStudy (mokeira):

lol..i was actually kinda stuck. @satellite73

OpenStudy (anonymous):

so far, \[-4\pm \sqrt{16+48}/2a\]

OpenStudy (anonymous):

what do i do next? @satellite73

OpenStudy (mokeira):

a=3 be sure to substitute that too

OpenStudy (anonymous):

that would be 6.... so, what now?

OpenStudy (mokeira):

\[-4\pm \frac{ \sqrt{64} }{ 6 }\] find square root of 64